If P(x) =√x, show that P(x+h) - P(x) = h(√x+h + √x).

1
Expert's answer
2022-06-06T14:28:50-0400

"P(x)=\sqrt{x},"

then

"P(x+h)=\sqrt{x+h}."

Let's subtract and simplify:

"P(x+h) - P(x)=\sqrt{x+h}-\sqrt{x}="


"= \frac{(\sqrt{x+h}-\sqrt{x})(\sqrt{x+h}+\sqrt{x})}{\sqrt{x+h}+\sqrt{x}}="


"= \frac{(\sqrt{x+h})^2-(\sqrt{x})^2}{\sqrt{x+h}+\sqrt{x}}= \frac{(x+h)-x}{\sqrt{x+h}+\sqrt{x}}= \frac{h}{\sqrt{x+h}+\sqrt{x}}."


So, we showed that

"P(x+h) - P(x)= \frac{h}{\sqrt{x+h}+\sqrt{x}}."